# DeSTA2.5-Audio: Toward General-Purpose Large Audio Language Model with Self-Generated Cross-Modal Alignment

### v1

@@ IV-A Dataset
All responses were generated using the vLLM toolkit [112], with decoding parameters set to a temperature of 0.05 and a top-p value of 1.0. This process yielded a large-scale dataset of approximately 5 million audio–prompt–response triplets, which we refer to as DeSTA-AQA5M. This dataset serves as the training corpus for DeSTA2.5-Audio. In this work, while we primarily study the performance of Llama3.1-8B-Instruct  [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct [4], the same data construction process can be easily generated by other LLMs within one day on a single NVIDIA A100-80GB GPU.
@@ IV-B Model Specification and Training Setup
We adopt Llama3.1-8B-Instruct and Whisper-large-v3444openai/whisper-large-v3 as the foundational components. We employ a six-layer Q-former architecture [24] with 64 queries as the modality adapter. The query vectors attend to intermediate hidden states from Whisper encoder layers 8, 16, 24, and 32, allowing the model to capture multi-scale acoustic features [113, 114]. For the optional linguistic representation, we provide offline transcriptions for datasets originating from the speech domain. For datasets in the audio and music domains, no transcription is used, and only continuous embeddings derived from the Q-Former are utilized during training. During inference, we employ a lightweight pre-trained voice activity detection (VAD) model [115] to identify the presence of human speech in audio inputs and conditionally activate the Whisper decoder when necessary.
The model implementation is based on the Transformers library [116]. It consists of 8.8 billion total parameters, with 131 million trainable parameters. Training is performed for five epochs using the Adam optimizer [117], a cosine annealing learning rate schedule, and 2000 warm-up steps. The training is conducted on a cluster of 8 NVIDIA A100-80GB GPUs, with a global batch size of 96 and an initial learning rate of 1e-4. The total number of training steps is approximately 250,000.
@@ IV-C Comparison Study
To assess the robustness and effectiveness of our proposed framework, we construct a controlled subset of 500,000 samples from DeSTA-AQA5M. This subset enables systematic analysis of how various training target generation alternatives and model configurations affect model performance. Unless otherwise specified, all comparison experiments are trained for ten epochs using identical model architectures and hyperparameter settings.
In the dataset-level comparison, we fix both the audio descriptions and prompts, and vary only the LLM responsible for generating the training targets. We consider Llama3.1-8B-Instruct [4], Qwen2.5-7B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct [6], Gemma3-12B-it [^6] https://huggingface.co/google/gemma-3-12b-it [7], and Llama3.1-70B-Instruct [^7] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct [4]. This setup facilitates an analysis of how variations in the output distribution across different LLMs influence downstream training outcomes under otherwise controlled conditions. Additionally, to isolate the impact of prompt diversity, we construct a caption-style dataset using a single fixed prompt (i.e., “What can you hear from the audio?”) as proposed in our prior work [9]. For the model-level comparison, we augment the LLM backbone with LoRA adapters [37] of rank 8, applied to the query, key, and value projections of each transformer layer. This increases the number of trainable parameters to 140 million. Additionally, we include a five-epoch variant in which only the number of updates and the learning rate scheduler differ from the ten-epoch setting.
@@ V-A Benchmarks
SAKURA [29] is a benchmark developed to evaluate single-hop and multi-hop reasoning in LALMs, referred to as SAKURA-Single and SAKURA-Multi, respectively. The benchmark consists of multiple-choice questions. Specifically, single-hop questions assess basic auditory perception, such as identifying sounds like “What animal is making this sound?” In contrast, multi-hop questions require combining auditory cues with external world knowledge and reasoning beyond the immediate input. For example, a multi-hop question might ask, “Is the sound source a mammal or not?”, which requires both recognizing the sound and knowing which animals are classified as mammals. These benchmarks are particularly useful for assessing the internal knowledge of LALM.
@@ V-B Cascade Baselines
V-C Model Inference Setup
VI Main Results
TABLE I: Evaluation results across Dynamic-SUPERB Phase-1, MMAU, SAKURA and Speech-IFEval. $\dagger$ denotes the cascade reference system result from [36].
@@ VI-A Results on Dynamic-SUPERB Phase-1, MMAU, SAKURA and Speech-IFEval
Table I presents results across benchmarks, compared to several representative models. To contextualize the benchmark performance, we first consider tasks related to auditory perception, including Dynamic-SUPERB Phase-1 and MMAU, which evaluate models on audio understanding. The SAKURA benchmark, which spans both single-hop and multi-hop reasoning, serves as a bridge between basic perception and higher-level reasoning. Finally, we assess general knowledge and instruction-following through benchmarks like SAKURA-Multi and Speech-IFEval.
@@ VI-A1 Overall Trend
Compared to DeSTA2, which focused primarily on speech data, DeSTA2.5-Audio shows improved performance attributed to the broader coverage of environmental and musical audios in addition to speech. This demonstrates that our proposed data construction strategy is robust and can be effectively extended to other domains. Interestingly, while Qwen2-Audio-Instruct leveraged a massive 510,000 hours of training data, DeSTA2.5-Audio achieves competitive or superior performance using only 7,000 hours. This highlights the efficiency and effectiveness of our training methodology, which enables strong results with significantly less data.
@@ VI-A2 Task Performance
DeSTA2.5-Audio consistently achieves best performance across a variety subdomains. Within speech-related domains, the model demonstrates strong capabilities not only in content and semantic comprehension but also in more advanced perception, such as paralinguistics, degraded speech, and speaker identity in the Dynamic-SUPERB phase-1. Beyond speech, DeSTA2.5-Audio also performs competitively in the audio and music tracks of the MMAU benchmark, underscoring its versatility in processing diverse auditory inputs.
In the single-hop setting (SAKURA-Single), which requires grounding questions to specific audio segments, Qwen2-Audio-Instruct outperforms DeSTA2.5-Audio (81.20 vs. 76.65). This is likely attributable to Qwen2-Audio-Instruct’s extensive instruction tuning, which enhances its ability to respond to direct queries. However, the performance dynamics shift dramatically in the multi-hop setting (SAKURA-Multi). Here, DeSTA2.5-Audio emerges as the top performer with a score of 69.85, while other models suffer significant drops. For example, Qwen2-Audio-Instruct declines by over 30 points (from 81.20 to 49.10), and LTU-AS drops from 40.90 to 18.10, underscoring their struggles with compositional reasoning. A similar trend can be observed in the Speech-IFEval benchmark. For instance, models such as LTU-AS and SALMONN exhibit low IFrate and high forgetting rates, indicating a significant degradation in both text-based knowledge and instruction-following capabilities. In contrast, DeSTA2.5-Audio and DeSTA2 demonstrate remarkable robustness on the Speech-IFEval benchmark. Their performance remains comparable to or even surpasses the text-based reference systems [36]. The results show that our proposed approach yields a significant advantage in multi-hop reasoning and instruction-following performance, while also delivering strong results on core auditory perception tasks.
Figure 3: LALM performance on the Dynamic-SUPERB Phase-2 benchmark [33], relative to the ASR+LLM baseline. Each cell shows a model’s relative score in a specific domain, with darker blue indicating a higher performance rank. The bottom rows summarize each model’s domain win count (positive scores) and average performance across all domains.
TABLE II: Evaluation results on VoiceBench. Performance is measured using task-specific metrics, where higher scores indicate better performance.
@@ VI-B Results on Dynamic-SUPERB Phase-2
A closer analysis reveals that in certain domains, such as “Quality Assessment, Speech, Voice, Hearing Disorder”, and “Phonetics, Phonology, and Prosody”, our model underperforms the cascade baseline, as indicated by negative relative scores. This suggests that the audio attributes related to addressing these types of tasks are currently underrepresented in our training corpus, thereby limiting generalization performance in these areas. However, the relative scores remain relative to the ASR+LLM baseline, implying marginal deviation despite underperformance. In contrast, some systems exhibit severe performance degradation, for example, LTU-AS in Phonetics, Phonology, and Prosody, and SALMONN in Signal Characteristics Analysis.
@@
TABLE III: Evaluation results across various configurations on representative benchmarks. “Backbone” and “Dataset” refer to the LLM used for LALM training and for dataset generation, respectively. The label “(1-P)” indicates that the dataset was generated using a single descriptive prompt[9]. “PPL” denotes the perplexity of the training targets as computed by the corresponding backbone LLM. “N/A” indicates that the result is unavailable due to degeneration.
@@
With 5 epochs
D1 Llama3.1 Llama3.1 1.61 90.77 70.42 57.50 58.08 53.60 66.56 55.26 57.66 48.20 53.70 70.25 58.80 92.94 -0.62
D2 Llama3.1 Llama3.1 (1-P) 1.58 77.55 60.58 50.57 42.00 40.00 53.51 54.95 54.05 40.12 49.70 58.15 51.70 90.21 -3.54
D3 Llama3.1 Qwen2.5 3.54 74.82 62.67 53.21 53.45 57.70 59.91 57.06 55.56 43.41 52.00 44.30 53.40 91.80 -1.85
VI-C Results on VoiceBench
Table II presents the results on VoiceBench, where the audio input consists of spoken instructions, such as questions or commands, and the model is expected to generate appropriate responses. Unlike previous benchmarks that focus mainly on auditory understanding or instruction-following, VoiceBench emphasizes general knowledge and interactive behavior. To align with these goals, we guide DeSTA2.5-Audio using a system prompt that encourages more natural and conversational responses.
In terms of overall performance, the ASR+LLM pipeline proves to be a very strong baseline, achieving an overall score of 77.48 and outperforming most fusion-based approaches. This indicates that the tasks in VoiceBench are primarily content-driven and can be effectively addressed by accurate transcription paired with a powerful LLM. For comparison, OpenAI’s proprietary GPT-4o-Audio achieves the highest score of 86.75 on VoiceBench, establishing the current state-of-the-art in the field. Despite the simple of cascade systems, most open-source fusion models are still unable to exceed this baseline. This highlights a notable research gap in the field of recent LALM development. These models differ considerably in how they represent audio input. For example, Moshi [47] and GLM-4-Voice [48] adopt unit-based representations, while DiVA [21], Qwen2-Audio-Instruct [17], Phi-4 [18], VITA-1.5 [46] use continuous representations. Furthermore, several models such as Moshi, GLM-4-Voice, and VITA-1.5 are specifically designed for speech-to-speech interaction, making their training procedure more difficult than other speech-to-text model. Among all fusion-based models evaluated, DeSTA2.5-Audio achieves an overall score of 74.52, outperforming models such as VITA-1.5 (64.53) and Qwen2-Audio-Instruct (55.80). It demonstrates balanced and consistent results across a broad range of subtasks, including AlpacaEval, CommonEval, and OpenbookQA. This performance is largely attributed to the strong language ability of DeSTA2.5-Audio, where it not only follows the system prompt to adapt its behavior into an interactive style, but also responds appropriately based on its inherent knowledge.

### v2

@@ IV-A Dataset
All responses were generated using the vLLM toolkit [63], with decoding parameters set to a temperature of 0.05 and a top-p value of 1.0. This process yielded a large-scale dataset of approximately 5 million audio–prompt–response triplets, which we refer to as DeSTA-AQA5M. This dataset serves as the training corpus for DeSTA2.5-Audio. In this work, while we primarily study the performance of Llama3.1-8B-Instruct [24], the same data construction process can be easily generated by other LLMs within one day on a single NVIDIA A100-80GB GPU.
@@ IV-B Model Specification and Training Setup
We adopt Llama3.1-8B-Instruct and Whisper-large-v3 as the foundational components. We employ a six-layer Q-former architecture [66] with 64 queries as the modality adapter. The query vectors attend to intermediate hidden states from Whisper encoder layers 8, 16, 24, and 32, allowing the model to capture multi-scale acoustic features [121, 120]. For the optional linguistic representation, we provide offline transcriptions for datasets originating from the speech domain. For datasets in the audio and music domains, no transcription is used, and only continuous embeddings derived from the Q-Former are utilized during training. During inference, we employ a lightweight pre-trained voice activity detection (VAD) model [96] to identify the presence of human speech in audio inputs and conditionally activate the Whisper decoder when necessary.
The model implementation is based on the Transformers library [109]. It consists of 8.8 billion total parameters, with 131 million trainable parameters. Training is performed for five epochs using the Adam optimizer [58], a cosine annealing learning rate schedule, and 2000 warm-up steps. The training is conducted on a cluster of 8 NVIDIA A100-80GB GPUs for five days, with a global batch size of 96 and an initial learning rate of 1e-4. The total number of training steps is approximately 250,000.
@@ IV-C Comparison Study
To assess the robustness and effectiveness of our proposed framework, we construct a controlled subset of 500,000 samples from DeSTA-AQA5M. This subset enables systematic analysis of how various training target generation alternatives and model configurations affect model performance. All comparison experiments are trained for ten epochs using identical model architectures and hyperparameter settings.
In the dataset-level comparison, we fix both the audio descriptions and prompts, and vary only the LLM responsible for generating the training targets. We consider Llama3.1-8B-Instruct [24], Qwen2.5-7B-Instruct [115], Gemma3-12B-it [95], and Llama3.1-70B-Instruct [24]. This setup facilitates an analysis of how variations in the output distribution across different LLMs influence downstream training outcomes under otherwise controlled conditions. Additionally, to isolate the impact of prompt diversity, we construct a caption-style dataset using a single fixed prompt (i.e., “What can you hear from the audio?”) as proposed in our prior work [73]. For the model-level comparison, we augment the LLM backbone with LoRA adapters [45] of rank 8, applied to the query, key, and value projections of each transformer layer. This increases the number of trainable parameters to 140 million.
@@ V-A Benchmarks
SAKURA [117] is a benchmark developed to evaluate single-hop and multi-hop reasoning in LALMs, referred to as SAKURA-Single and SAKURA-Multi, respectively. The benchmark consists of multiple-choice questions that include gender, emotion, animal and language classification tasks. Specifically, single-hop questions assess basic auditory perception, such as identifying sounds like “What animal is making this sound?” In contrast, multi-hop questions require combining auditory cues with world knowledge and reasoning beyond the immediate input, such as “Is the sound source a mammal or not?”. These benchmarks are particularly useful for assessing the internal knowledge of LALM.
@@ V-B Cascade Baselines
TABLE II: Evaluation results across Dynamic-SUPERB Phase-1, MMAU, SAKURA and Speech-IFEval. $\dagger$ signifies the cascade reference system result from [74], and $\uparrow$ denotes higher scores are better.
@@
V-C Model Inference Setup
(a) Representation Extraction
(b) Similarity-based Classification
Figure 3: Illustration of the CLAP-style evaluation using the backbone LLM for embedding extraction. (a) Representation Extraction: Inputs are fed into the LLM. We utilize the hidden state of the last token as the aggregated representation for the input. (b) Similarity-based Classification: The derived audio representation is compared against the text representations of all candidate class labels using cosine similarity. The label with the highest similarity score (in this example, “Male”) is selected as the prediction.
V-D Representation-Level Evaluation Setup
Beyond standard instruction-based evaluation, we adopt a CLAP-style protocol using similarity-based classification [25, 112] to assess the representation-level cross-modal alignment quality, as illustrated in Figure 3. Since our autoregressive models do not yield a single explicit embedding vector for variable-length inputs, we utilize the backbone LLM as a universal encoder [55, 46]. As depicted in Figure 3(a), we feed either the audio (via the modality adapter) or the text sequence into the LLM. We then append a summarization prompt (e.g., “summarize the audio in one word:”) to the input and extract the hidden state of the last token as the aggregated representation [54].
We conduct this analysis on SAKURA-Single, targeting gender (2 classes), emotion (5 classes), and animal sounds (9 classes). Class labels are converted into text embeddings using descriptive templates (e.g., “The speaker is a male”, “The speaker feels happy” and “There is a cat”). As depicted in Figure 3(b), classification is performed by identifying the text representation with the highest cosine similarity to the audio embedding. Unless otherwise specified, we report scores from the final transformer layer. Higher accuracy suggests that the LLM interprets the audio representations similar to the semantic concept.
VI Main Results
@@ VI-A Results on Dynamic-SUPERB Phase-1, MMAU, SAKURA and Speech-IFEval
Table II presents results across benchmarks, compared to several representative models. To contextualize the benchmark performance, we first consider tasks related to auditory perception, including Dynamic-SUPERB Phase-1 and MMAU, which evaluate models on audio understanding. The SAKURA benchmark, which spans both single-hop and multi-hop reasoning, serves as a bridge between basic perception and higher-level reasoning. Finally, we assess general knowledge and instruction-following through benchmarks like SAKURA-Multi and Speech-IFEval.
Figure 4: LALM performance on the Dynamic-SUPERB Phase-2 benchmark [48], relative to the ASR+LLM baseline. Each cell shows a model’s relative score in a specific domain, with darker blue indicating a higher performance rank. The bottom rows summarize each model’s domain win count (i.e., positive scores) and average performance across all domains.
@@ VI-A1 Overall Trend
Compared to DeSTA2, which focused primarily on speech data, DeSTA2.5-Audio shows improved performance attributed to the broader coverage of environmental sounds and musical audios in addition to speech. This demonstrates that our proposed self-generated data construction strategy is general and can be effectively extended to other domains. Interestingly, while Qwen2-Audio-Instruct leveraged a massive 510,000 hours of training data, DeSTA2.5-Audio achieves competitive or superior performance using only 7,000 hours. This highlights the efficiency and effectiveness of our training methodology, which enables strong results with significantly less data.
@@ VI-A2 Task Performance
DeSTA2.5-Audio consistently achieves the best performance across a variety of subdomains. Within speech-related domains, the model demonstrates strong capabilities not only in content and semantic comprehension but also in more advanced perception, such as paralinguistics, degraded speech, and speaker identity in the Dynamic-SUPERB phase-1. Beyond speech, DeSTA2.5-Audio also performs competitively in the audio and music tracks of the MMAU benchmark, underscoring its versatility in processing diverse auditory inputs.
In the single-hop setting (SAKURA-Single), which requires grounding questions to specific audio segments, Qwen2-Audio-Instruct outperforms DeSTA2.5-Audio (81.20 vs. 76.65). This is likely attributable to Qwen2-Audio-Instruct’s extensive instruction tuning, which enhances its ability to respond to direct queries. However, the performance dynamics shift dramatically in the multi-hop setting (SAKURA-Multi). Here, DeSTA2.5-Audio emerges as the top performer with a score of 69.85, while other models suffer significant drops. For example, Qwen2-Audio-Instruct declines by over 30 points (from 81.20 to 49.10), and LTU-AS drops from 40.90 to 18.10, underscoring their struggles with compositional reasoning. A similar trend can be observed in the Speech-IFEval benchmark. For instance, models such as LTU-AS and SALMONN exhibit low IFrate and high forgetting rates, indicating a significant degradation in both text-based knowledge and instruction-following capabilities. In contrast, DeSTA2.5-Audio and DeSTA2 demonstrate remarkable robustness on the Speech-IFEval benchmark. Their performance remains comparable to or even surpasses the text-based reference systems [74]. These benchmarks demand extensive textual knowledge and complex instruction-following, effectively serving as out-of-domain tests for the LALMs. The results show that our proposed approach yields a significant advantage in multi-hop reasoning and instruction-following performance, while also delivering strong results on core auditory perception tasks.
TABLE III: Evaluation results on VoiceBench. Performance is measured using task-specific metrics, where higher scores indicate better performance.
@@ VI-B Results on Dynamic-SUPERB Phase-2
A closer analysis reveals that in certain domains, such as “Quality Assessment, Speech, Voice, Hearing Disorder”, and “Phonetics, Phonology, and Prosody”, our model underperforms the cascade baseline, as indicated by negative relative scores. This suggests that the audio attributes required for these tasks are currently underrepresented in our training corpus, thereby limiting generalization performance in these areas. However, the relative scores remain relative to the ASR+LLM baseline, implying marginal deviation despite underperformance. In contrast, some systems exhibit severe performance degradation, for example, LTU-AS in Phonetics, Phonology, and Prosody, and SALMONN in Signal Characteristics Analysis.
@@
VI-C Results on VoiceBench
Table III presents the results on VoiceBench, where the audio input consists of spoken instructions, such as questions or commands, and the model is expected to generate appropriate responses. Unlike previous benchmarks that focus mainly on auditory understanding or instruction-following, VoiceBench emphasizes general knowledge and interactive behavior. To align with these goals, we guide DeSTA2.5-Audio using a system prompt that encourages more natural and conversational responses.
In terms of overall performance, the ASR+LLM pipeline proves to be a very strong baseline, achieving an overall score of 77.48 and outperforming most fusion-based approaches. This indicates that the tasks in VoiceBench are primarily content-driven and can be effectively addressed by accurate transcription paired with a powerful LLM. For comparison, OpenAI’s proprietary GPT-4o-Audio achieves the highest score of 86.75 on VoiceBench, establishing the current state-of-the-art in the field. Despite the simplicity of cascade systems, most open-source fusion models are still unable to exceed this baseline. This highlights a notable research gap in the field of recent LALM development. These models differ considerably in how they represent audio input. For example, Moshi [22] and GLM-4-Voice [125] adopt unit-based representations, while DiVA [42], Qwen2-Audio-Instruct [17], Phi-4 [1], VITA-1.5 [30] use continuous representations. Furthermore, several models such as Moshi, GLM-4-Voice, and VITA-1.5 are specifically designed for speech-to-speech interaction, making their training procedure more difficult than other speech-to-text models. Among all fusion-based models evaluated, DeSTA2.5-Audio achieves an overall score of 74.52, outperforming models such as VITA-1.5 (64.53) and Qwen2-Audio-Instruct (55.80). It demonstrates balanced and consistent results across a broad range of subtasks, including AlpacaEval, CommonEval, and OpenbookQA. This performance is largely attributed to the strong language ability of DeSTA2.5-Audio, where it not only follows the system prompt to adapt its behavior into an interactive style, but also responds appropriately based on its inherent knowledge.
VI-D Representation-Level Analysis
TABLE IV: Comparison of instruction-based QA and similarity-based classification on SAKURA-Single. Higher value indicates better cross-modal alignment quality.
Method Gender Emotion Animal
# classes 2 5 9
Random 50.0 20.0 11.1
Instruction-based QA 89.2 62.2 58.8
Similarity-based classification 95.2 56.2 66.2
Beyond standard instruction-following evaluations, we investigate the cross-modal alignment quality between audio and text representations within DeSTA2.5-Audio, as described in Section V-D. Table IV compares standard text generation (“Instruction-based QA”) with implicit representation alignment (“Similarity-based classification”). The results demonstrate that DeSTA2.5-Audio establishes robust cross-modal grounding and encodes highly discriminative audio features, showing performance comparable to instruction-based evaluation. This indicates that DeSTA2.5-Audio successfully integrates audio representations into the backbone LLM’s semantic space. Furthermore, our results reveal an interesting trade-off. For purely acoustic tasks like Gender and Animal sounds, the model’s internal representations outperform text generation. This suggests that some audio information may be slightly diluted or lost during the instruction-following and text decoding stages. However, the opposite occurs for Emotion. Since emotion depends not just on acoustic features but also on spoken content, the representation-based approach may miss semantic nuances, whereas the instruction-based approach can leverage them for more accurate emotion recognition.
Figure 5: Epoch-wise evolution of cross-modal alignment. The curves depict similarity-based classification accuracy during training.
Figure 6: Layer-wise analysis of cross-modal alignment within the LLM. We measure similarity-based classification accuracy using hidden states from each transformer layer. The sharp rise at the deeper layers indicates that audio-text integration occurs through a deferred process within the semantic space.
Figure 5 illustrates the evolution of alignment quality across training epochs. We observe that simple acoustic attributes, such as gender, are learned rapidly, achieving near-convergence by the first epoch and more complex concepts like emotion and animal sounds demonstrate a steady upward trend as training progresses. This trajectory confirms that DeSTA2.5-Audio effectively drives the model to progressively refine its understanding of multifaceted auditory concepts over time.
Figure 6 presents a detailed analysis of the accuracy based on hidden states extracted from different layers of the backbone LLM. Interestingly, a distinct phase transition phenomenon is observed. Accuracy remains near random levels in the shallow layers, but exhibits a sharp surge around layer 22. This finding implies that the model does not enforce immediate audio-text alignment in the input layers. Instead, the end-to-end optimization allows the raw acoustic features to be progressively integrated into the LLM’s semantic space in the deeper layers. This pattern is consistently observed in the comparative experiments in Section VII as well. We hope these findings inspire broader investigation into the underlying dynamics of cross-modal learning, fostering a deeper understanding of interpretability in multimodal foundation models.
@@ VI-D Representation-Level Analysis
TABLE V: Evaluation results across various configurations on representative benchmarks. “Backbone” and “Dataset” refer to the LLM used for LALM training and for dataset generation, respectively. The label “(1-P)” indicates that the dataset was generated using a single descriptive prompt[73]. “PPL” denotes the perplexity of the training targets as computed by the corresponding backbone LLM. “N/A” indicates that the result is unavailable due to degeneration.